OpenStudy (anonymous):

what is the value of x in the quadratic formula 3(x-2)^2+7=0 help me please

OpenStudy (shubhamsrg):

you comfortable with complex nos. ?

OpenStudy (anonymous):

hmm. anything :)

OpenStudy (shubhamsrg):

let (x-2)^2 = y => 3y +7 =0 whats the value of y from here?

OpenStudy (anonymous):

there is no y. the formulan is the X= -b + and - *square root* b^2-4ac over 2a

OpenStudy (shubhamsrg):

you answer my question and i'll ans your query then..

OpenStudy (anonymous):

what do you mean? do you know the answer?

OpenStudy (angela210793):

 but if u want complex solutions use